Find the account balance if the account starts with Php 100, has an annual rate of 4%, and the money left in the account for 12 years.
Answer by nerdybill(7384)

Find the account balance if the account starts with Php 100, has an annual rate of 4%, and the money left in the account for 12 years.
.
Assuming simple interest:
A = P(1+r)^t
where
P is 100
r is .04
t is 12
.
A = P(1+r)^t
A = 100(1+.04)^12
A = 100(1.04)^12
A = 100(1.601032218567680790102016)
A = Php 160.10